# Recent Advances in OOD Detection: Problems and Approaches

### v2

@@ 5.2. Few-shot OOD Detection Approaches
Studies. The direct way of adapting a large pre-trained model with several ID samples is fine-tuning.  Ming and Li (2024) and  Dong et al. (2023) conduct studies on the impact of fine-tuning on OOD detection within the context of VLMs, with a special focus on CLIP.  Ming and Li (2024) pay more attention to the role of parameter-efficient fine-tuning (PEFT) methods and the OOD score. The MCM score, introduced by  Ming et al. (2022b) as an innovative measure, alongside methods based on prompt learning for detecting OOD instances, is recognized for its effectiveness. Similarly,  Dong et al. (2023) conduct an extensive comparison between various fine-tuning approaches, such as PEFT and traditional methods, and discover that PEFT exhibits superior performance in detecting OOD instances, which echos with the conclusion of  ming2023does. Following this,  Kim et al. (2024) believe that Finetune Like You Pretrain (FLYP), a fine-tuning method, deserves attention due to its good performance on classification tasks. Specifically, FLYP is to mimic contrastive language-image pre-training as CLIP. After comparing the performance of FLYP and PEFT methods on zero-shot OOD detection,  Kim et al. (2024) find FLYP yields better OOD detection performance than PEFT. It should be noted that while  Fort et al. (2021) discuss “few-shot” OOD detection, their approach relies on a small number of outlier examples rather than ID samples, making it unsuitable for inclusion in this section.
Fine-tuning based. A prevalent method for fine-tuning CLIP using limited ID data involves prompt learning (Zhou et al., 2022b, 2023), where the fine-tuning focuses on the prompt’s contextual words, maintaining the pre-trained parameters unchanged. To distinguish between ID and OOD data, both LoCoOp  (Miyai et al., 2023a) and ID-like Prompting Learning  (Bai et al., 2024) are designed to enhance the learning of context vectors from the near-ID feature perspective. LoCoOp employs entropy maximization to distance ID-irrelevant local features (such as the backgrounds in ID images) from the textual embeddings of ID classes. Similarly, ID-like Prompting Learning generates ID-like data—outliers within the vicinity of ID samples—to refine the context vector. Additionally, it incorporates a diversity loss to enhance the variance among the sampled OOD data. However, recent work EOK-Prompt (Zeng et al., 2024) and GalLop (Lafon et al., 2024) argues that optimizing only a prompt using numerous features from an image wastes valuable information. Instead, EOK proposes an orthogonal method to optimize the novel local prompt, making better use of the limited image data, while GalLop optimizes global prompts and local prompts by leveraging global and local features, respectively. Furthermore, based on the mentioned studies, DSFG  (Dong et al., 2023) posits that traditional fine-tuning approaches applied to CLIP may inadvertently lead to a loss of critical OOD knowledge. To address this issue, DSFG adopts a strategy to retain broad OOD knowledge by merging the original features with those modified through fine-tuning, followed by the training of a classifier. DSFG’s plug-and-play capability makes it seamlessly compatible with all fine-tuning-based approaches, enhancing its practicality and value. The aforementioned few-shot tuning methods inevitably suffer from overfitting on limited-shot instances. Chen et al. (2024b) proposed a training-free few-shot OOD detection method, Dual-Adapter(Chen et al., 2024b), which constructs two types of adapters by extracting positive and negative features to aid in OOD detection. The currently mentioned few-shot methods lack OOD supervision, so CLIP-OS (Sun et al., 2024) tries to find such supervision and also achieves stunning results
@@ 5.3. Full-shot OOD Detection Approaches
Overview. This setup is generally less realistic than the first two (zero-shot and few-shot). However, we list them separately to ensure a comprehensive review of existing methods across the spectrum. Given the full set of ID data and corresponding labels, VLMs can enhance OOD detection significantly by fine-tuning. Moreover, a novel task called “PT-OOD” detection is introduced.
@@ 6.2. Experimental Protocols
Subsequently, OOD data is categorized into near-OOD and far-OOD based on the degree of covariate shift from ID data. This categorization corresponds to dividing OOD detection tasks into near-OOD and far-OOD detection. It is evident the near-OOD detection task is more challenging, however, numerous methods  (Fort et al., 2021; Ming et al., 2022b; Bai et al., 2024) have demonstrated excellent performance in this area.
Recently,  Yang et al. (2023b); Bai et al. (2023) propose that we should consider cases where covariate shift occurs in ID data, which is not taken into account previously. This is crucial to prevent the loss of model generalization. The samples mentioned earlier are termed as cs-ID data, an abbreviation for “covariate shift ID” data. Consequently, a new experimental protocol has been explored, called full-spectrum OOD detection. During the testing phase, the model is expected to identify near-OOD and far-OOD instances. Additionally, it should refuse to provide predictions for the OOD data and accurately predict ID and cs-ID data.
@@ 6.3. Application
Table 3. Summary of Datasets. The CARLA System is a simulation platform designed for evaluating OOD Detection in the field of autonomous driving, hence its entire row is filled with dashes. Other “-” symbols represent numbers that vary depending on usage. More detailed descriptions can be found in the code repo.
TASK Dataset Name Data Type # Classes # Samples Papers
Image Classification CIFAR-10 Images 10 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022)
CIFAR-100 Images 100 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022)
MNIST Images 10 70,000 (Hendrycks and Gimpel, 2017; Zhao et al., 2024)
ImageNet-1K Images 1,000 1,431,167 (Ming et al., 2022b; Wang et al., 2023c)
iNaturalist Images 5,089 675,170 (Ming et al., 2022b; Wang et al., 2023c)
SUN Images 397 108,754 (Ming et al., 2022b; Wang et al., 2023c)
Places Images >205 >2,500,000 (Ming et al., 2022b; Wang et al., 2023c)
Textures Images 47 5,640 (Ming et al., 2022b; Wang et al., 2023c)
Semantic Segmentation Cityscapes Images - 25,000 (Gao et al., 2023b)
Road Anomaly Dataset Images - 100 (Gao et al., 2023b)
Object Detection PASCAL VOC Images 20 2,913 (Nguyen et al., 2024; Du et al., 2022b)
BBD100K Video->Image Variable 100,000 (Nguyen et al., 2024; Du et al., 2022b)
Autonomous Driving CARLA System - - - (Mao et al., 2024)
Medical Image Analysis Kvasir-Capsul Images 4,741,621 (Tan et al., 2024)
Text Category News Category Text - 210,000 (Arora et al., 2021)
SST-2 Text - 215,154 (Arora et al., 2021)
Intent Detection CLINC150 Text 150 22,500 (Zhan et al., 2022)
Banking Text 77 13,083 (Zhan et al., 2022)
StackOverflow Text 20 20,000 (Zhan et al., 2022)
Audio MSCW Audio - >23400000 (Bukhsh and Saeed, 2023)
Vocalsound Audio - 21,024 (Bukhsh and Saeed, 2023)
Graph data TU Graph data - Variable (Wang et al., 2024)
OGB Graph data - Variable (Wang et al., 2024)
@@ 6.3.1. Computer Vision
  • Image Classification. The majority of tasks discussed in this paper focus on OOD detection within the realm of image classification. In such task scenarios, commonly utilized ID datasets include MNIST  (Xiao et al., 2017), CIFAR-10  (Krizhevsky et al., 2009), CIFAR-100  (Krizhevsky et al., 2009), and ImageNet-1K  (Deng et al., 2009). Various OOD datasets are constructed accordingly for evaluating different methods  (Yang et al., 2022b), with the most frequently employed datasets being iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places (Zhou et al., 2017) and Textures (Cimpoi et al., 2014). Additionally, full-spectrum OOD is typically evaluated on three benchmarks: DIGITS, OBJECTS, and COVID, as proposed by  Yang et al. (2023b).
@@
  • Semantic Segmentation. Recent works  (Gao et al., 2023b) have started delving into the dense OOD detection task, also known as anomaly segmentation. The datasets used for evaluation include the Cityscapes dataset  (Cordts et al., 2016), the Road Anomaly dataset  (Lis et al., 2019), and the recently developed SOOD-ImageNet  (Bacchin et al., 2024).
  • Object Detection. The application of methods related to OOD detection in the field of object detection is relatively nascent, with only a few studies exploring this area  (Du et al., 2022b). Evaluations are commonly performed using datasets like PASCAL-VOC  (Everingham et al., 2015) and Berkeley DeepDrive-100K  (Yu et al., 2018).
  • Autonomous Driving. Autonomous driving has long been a crucial practical application of OOD Detection. Recently, Mao et al. (2024) utilized the CARLA (Dosovitskiy et al., 2017) system to simulate and evaluate the performance of OOD Detection in autonomous driving scenarios.
  • Medical Image Analysis. In the field of medical image analysis, OOD detection is crucial. Depending on the specific category of medical image, OOD detection employs various datasets including CIFAR-10 and Kvasir-Capsul (Smedsrud et al., 2021).
OOD detection has significant applications across various fields such as human action recognition (Sim et al., 2023) and solar image analysis (Giger and Csillaghy, 2024; Das et al., 2024). For further details, please refer to the accompanying code repository due to space limitations. The same applies to the subsequent paragraphs.

### v3

@@ 5.2. Few-shot OOD Detection Approaches
Studies. The direct way of adapting a large pre-trained model with several ID samples is fine-tuning.  Ming and Li (2024) and  Dong et al. (2023) conduct studies on the impact of fine-tuning on OOD detection within the context of VLMs, with a special focus on CLIP.  Ming and Li (2024) pay more attention to the role of parameter-efficient fine-tuning (PEFT) methods and the OOD score. The MCM score, introduced by  Ming et al. (2022b) as an innovative measure, alongside methods based on prompt learning for detecting OOD instances, is recognized for its effectiveness. Similarly,  Dong et al. (2023) conduct an extensive comparison between various fine-tuning approaches, such as PEFT and traditional methods, and discover that PEFT exhibits superior performance in detecting OOD instances, which echos with the conclusion of  Ming and Li (2024). Following this,  Kim et al. (2024) believe that Finetune Like You Pretrain (FLYP), a fine-tuning method, deserves attention due to its good performance on classification tasks. Specifically, FLYP is to mimic contrastive language-image pre-training as CLIP. After comparing the performance of FLYP and PEFT methods on zero-shot OOD detection,  Kim et al. (2024) find FLYP yields better OOD detection performance than PEFT. It should be noted that while  Fort et al. (2021) discuss “few-shot” OOD detection, their approach relies on a small number of outlier examples rather than ID samples, making it unsuitable for inclusion in this section.
Few-shot OOD Detection. Few-shot OOD detection aims to leverage a limited amount of ID data to adapt LPMs, primarily through prompt learning or parameter-efficient fine-tuning. Early approaches such as CoOp (Zhou et al., 2022b) and ZegCLIP (Zhou et al., 2023) focus on optimizing the contextual words in prompts while keeping the backbone fixed, thus preserving the generalization ability of the pre-trained models. Building upon this, LoCoOp (Miyai et al., 2023a) and IDPL (Bai et al., 2024a) further enhance context vector learning from a near-ID perspective: LoCoOp maximizes entropy to separate ID-irrelevant local features (e.g., backgrounds) from textual embeddings of ID classes, while IDPL synthesizes ID-like outliers near the ID boundary and employs a diversity loss to encourage variance among sampled OOD candidates. In addition, MMFT (Kim and Hwang, 2025) projects image and text features into a shared hyperspherical space and introduces a cross-modal alignment loss function to promote the alignment of image and text representations in the hyperspherical space. During training, the loss function brings the image and text embeddings closer together while effectively distinguishing OOD texts from ID texts. However, these prompt-only methods may underutilize the rich information contained in image features. Recent advances such as GalLoP (Lafon et al., 2024) and Local-Prompt (Zeng et al., 2025) address this limitation by jointly optimizing global and local prompts. GalLoP leverages both global and region-level features via prompt dropout and region-level regularization to improve generalization and robustness, whereas Local-Prompt introduces orthogonal local prompt optimization to better exploit limited image data.
Beyond prompt learning, DSFG (Dong et al., 2023) highlights that conventional fine-tuning can erode critical OOD knowledge. To mitigate this, DSFG merges original and fine-tuned features before classifier training, thereby retaining broad OOD awareness and offering plug-and-play compatibility with other fine-tuning-based methods. Recently, SCT (Yu et al., 2024a) proposes a self-calibrated tuning framework that dynamically balances ID classification and OOD regularization by introducing modulation factors based on sample uncertainty, enabling more reliable adaptation with only a few ID samples. Complementary to this, GaCoOp (Tong et al., 2025) addresses the gradient conflict between ID classification and OOD regularization in prompt optimization. By decomposing and projecting gradients, GaCoOp aligns the optimization directions, thereby mitigating the adverse effects of background bias and improving few-shot OOD detection performance, especially when data is scarce.
Despite these advancements, overfitting remains a challenge due to the scarcity of labeled data. To tackle this, Dual-Adapter (Chen et al., 2024c) proposes a training-free few-shot method that constructs positive and negative adapters in a plug-and-play manner, requiring only a handful of ID samples. Meanwhile, to address the intrinsic modality gap in VLMs, SUPREME (Wang et al., 2025a) introduces multi-modal prototypes and alignment modules, while CLIP-OS (Sun et al., 2024) regularizes the ID/OOD boundary by synthesizing OOD data through mixing ID features, achieving strong performance without external OOD supervision. Additionally, NegPrompt (Li et al., 2024c) further refines the boundary by optimizing negative prompts, demonstrating the continued evolution and diversification of few-shot OOD detection strategies.
@@ 5.3. Full-shot OOD Detection Approaches
Overview. While this setup is generally less realistic than the first two (zero-shot and few-shot), we list them separately to ensure a comprehensive review of existing methods across the spectrum. Given the full set of ID data and corresponding labels, VLMs can enhance OOD detection significantly by fine-tuning. Moreover, a novel task called “PT-OOD” detection is introduced.
@@ 6. Evaluation and Application
Table 4. Summary of Datasets. The CARLA System is a simulation platform designed for evaluating OOD Detection in the field of autonomous driving, hence its entire row is filled with dashes. “*”indicates that this dataset is divided by category, with some data used as ID and other data as OOD. “/” symbol in the Usage column indicates that this dataset can be used as both ID and OOD data.
TASK Dataset Name Data Type # Classes # Samples Reference Usage
Image Classification CIFAR-10 Images 10 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022) ID/OOD
CIFAR-100 Images 100 60,000 (Wang et al., 2023c; Esmaeilpour et al., 2022) ID/OOD
MNIST Images 10 70,000 (Hendrycks and Gimpel, 2017; Zhao et al., 2024) ID/OOD
ImageNet-1K Images 1,000 1,431,167 (Ming et al., 2022b; Wang et al., 2023c) ID
iNaturalist Images 5,089 675,170 (Ming et al., 2022b; Wang et al., 2023c) OOD
SUN Images 397 108,754 (Ming et al., 2022b; Wang et al., 2023c) OOD
Places Images >205 >2,500,000 (Ming et al., 2022b; Wang et al., 2023c) OOD
Textures Images 47 5,640 (Ming et al., 2022b; Wang et al., 2023c) OOD
Semantic Segmentation Cityscapes Images 30 25,000 (Gao et al., 2023a) ID
Road Anomaly Dataset Images > 5 60 (Gao et al., 2023a) OOD
Object Detection PASCAL VOC Images 20 2,913 (Nguyen et al., 2024; Du et al., 2022b) ID
BDD100K Images 10 100,000 (Nguyen et al., 2024; Du et al., 2022b) ID
MS-COCO Images 80 > 930 (Nguyen et al., 2024; Du et al., 2022b) OOD
OpenImages Images 601 1761 (Nguyen et al., 2024; Du et al., 2022b) OOD
Autonomous Driving CARLA System - - - (Mao et al., 2024) -
Medical Image Analysis Kvasir-Capsule* Images 14 4,741,621 (Tan et al., 2024) ID/OOD
Text Category News Category* Text 41 210,000 (Arora et al., 2021) ID/OOD
SST-2 Text 2 215,154 (Arora et al., 2021) ID/OOD
Intent Detection CLINC150* Text 150 22,500 (Zhan et al., 2022) ID/OOD
Banking77* Text 77 13,083 (Zhan et al., 2022) ID/OOD
StackOverflow* Text 20 20,000 (Zhan et al., 2022) ID/OOD
Audio MSCW Audio 31 >23,400,000 (Bukhsh and Saeed, 2023) ID/OOD
Vocalsound Audio 6 21,024 (Bukhsh and Saeed, 2023) ID/OOD
Graph data TU/OGB* Graph data 10 19,766 (Wang et al., 2024a) ID/OOD
@@ 6.2. Experimental Protocols
Subsequently, OOD data is categorized into near-OOD and far-OOD based on the degree of covariate shift from ID data. This categorization corresponds to dividing OOD detection tasks into near-OOD and far-OOD detection. It is evident the near-OOD detection task is more challenging, however, numerous methods  (Fort et al., 2021; Ming et al., 2022b; Bai et al., 2024a) have demonstrated excellent performance in this area.
Recently,  Yang et al. (2023b); Bai et al. (2023); Yu et al. (2024b) propose that we should consider cases where covariate shift occurs in ID data, which is not taken into account previously. This is crucial to prevent the loss of model generalization. The samples mentioned earlier are termed as cs-ID data, an abbreviation for “covariate shift ID” data. Consequently, a new experimental protocol has been explored, called full-spectrum OOD detection. During the testing phase, the model is expected to identify near-OOD and far-OOD instances. Additionally, it should refuse to provide predictions for the OOD data and accurately predict ID and cs-ID data.
@@ 6.3. Application
This section presents a comprehensive review of both academic and real-world applications of OOD detection. To promote further research and enable standardized evaluation, commonly used benchmark datasets are systematically summarized in Table 4.
To provide a comprehensive overview of recent advances in OOD detection, we first organize our discussion based on fundamental technical domains (Computer Vision, Natural Language Processing, and Beyond) as these domains have distinct methodological characteristics and theoretical foundations that shape how OOD detection is approached. This domain-based categorization allows us to systematically analyze the core technical innovations and challenges specific to each field. We then extend our analysis to an application-oriented perspective, as shown in Table 6.3, which bridges the gap between theoretical advances and practical implementations, illustrating how these domain-specific techniques are adapted and deployed in real-world scenarios.
Table 5. Representative applications of OOD detection across five major domains, highlighting key use cases in industrial, medical, safety, scientific, and general-purpose settings.
Domain Task Application Description Reference
Industrial Applications Autonomous Driving Detecting unseen obstacles or traffic scenarios to ensure driving safety in autonomous vehicles. (tes, )
3D Object Detection Identifying out-of-distribution objects in 3D industrial environments, such as warehouses or robotics. (3D, ; Li et al., )
Object Detection Finding anomalous or unknown items in industrial production lines. (Syed et al., 2024; Chen et al., 2024d; Zhang et al., 2025b)
Earth Observation Imagery Detecting novel land-cover types or changes in satellite and remote sensing images. (Le Bellier and Audebert, 2024)
Edge Computing Real-time detection of abnormal data on edge devices to enhance responsiveness and security. (Li et al., 2025)
Time Series Analysis Monitoring equipment by identifying abnormal patterns for predictive maintenance. (web, )
Healthcare and Biology Medical Image Analysis Detecting unknown diseases or anomalies in medical images (e.g., X-ray, MRI). (Dis, ; Oh et al., 2024; Pokhrel et al., 2024; Smedsrud et al., 2021)
Biological Classification Identifying previously unseen species and enhancing biodiversity monitoring. (Arslan et al., 2024)
Security and Surveillance Facial Authentication Detecting spoofed or unregistered faces for enhanced security. (Kahya et al., 2024)
Facial Expression Recognition Identifying unknown or abnormal facial expressions in surveillance. (Zhang et al., 2024b)
Human Action Recognition Detecting anomalous or dangerous actions in monitored environments. (Sim et al., 2023; Mandal et al., 2019; Xu et al., 2024b)
Audio Recognition Recognizing abnormal sounds such as alarms or intrusions. (Mazumder et al., 2021; Gong et al., 2022)
Scientific Research Astronomical Imaging Discovering unknown celestial objects or phenomena in astronomical data. (Zhou et al., 2024)
Solar Image Analysis Detecting novel solar activities or anomalies in solar observation images. (Giger and Csillaghy, 2024; Das et al., 2024)
Graph Neural Networks Detecting anomalies in complex scientific data such as molecular structures or social networks. (Bazhenov et al., 2022; Wu et al., 2023b; Jiang et al., 2025)
Mathematical Reasoning Identifying out-of-distribution problem types in automated reasoning systems. (Wang et al., 2024b)
Spiking Neural Networks Recognizing abnormal neural signals in bio-inspired computing models. (Terres-Escudero et al., ; Orchard et al., 2015)
General-Purpose Applications Image Classification Detecting unknown categories in image classification systems. (Yang et al., 2023a, 2022; Sun et al., 2021)
Semantic Segmentation Identifying unknown regions or objects in pixel-level segmentation tasks. (Gao et al., 2023a; Lis et al., 2019)
Text Classification Detecting unseen topics or intents in text data. (Zhan et al., 2024)
Intent Detection Recognizing new user intents in intelligent assistants and chatbots. (Wang et al., 2024c; Zhan et al., 2022)
Question-Answering Systems Detecting questions that are out-of-scope or unanswerable by the system. (Zhang et al., 2024c)
Document Classification Detecting anomalies in documents containing multiple data modalities. (Constantinou et al., 2024)
Reinforcement Learning Detecting abnormal behavior or novel states in agent-environment interactions. (Chen et al., 2025b; Sedlmeier et al., 2019)
@@ 6.3.1. Computer Vision
  • Image Classification. The majority of OOD detection research targets image classification, utilizing ID datasets such as MNIST (Xiao et al., 2017), CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and ImageNet-1K (Deng et al., 2009). Common OOD datasets include iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places (Zhou et al., 2017), and Textures (Cimpoi et al., 2014). Full-spectrum OOD evaluation often employs the DIGITS, OBJECTS, and COVID benchmarks (Yang et al., 2023b). In addition to the aforementioned methods, most current research classifies OOD datasets based on the difficulty of distinguishing them from ID datasets. The most popular approach is to use ImageNet-1k as the ID dataset, with SSB-hard (Vaze et al., 2022) and NINCO (Bitterwolf et al., 2023) categorized as near-OOD (representing categories that are difficult to distinguish), and iNaturalist (Van Horn et al., 2018), Textures (Cimpoi et al., 2014), and OpenImage-O (Wang et al., 2022c) categorized as far-OOD (representing categories that are easier to distinguish
@@
  • Semantic Segmentation. Recent works  (Gao et al., 2023a) have started delving into the dense OOD detection task, also known as anomaly segmentation. The datasets used for evaluation include the Cityscapes dataset  (Cordts et al., 2016), the Road Anomaly dataset  (Lis et al., 2019), and the recently developed SOOD-ImageNet  (Bacchin et al., 2024).
  • Object Detection. OOD detection in object detection is relatively new (Du et al., 2022b), with evaluations on datasets like PASCAL-VOC (Everingham et al., 2015) and Berkeley DeepDrive-100K (Yu et al., 2018). Novel approaches include Situation Monitor, which uses diversity-driven budding ensemble architectures for zero-shot OOD detection in Transformer-based models (Syed et al., 2024), Proto-OOD, which enhances OOD object detection with prototype feature similarity (Chen et al., 2024d), and VisTa, which employs visual-contextual and text-augmented methods for zero-shot object-level OOD detection (Zhang et al., 2025b).
  • 3D Object Detection. OOD detection in LiDAR-based 3D object detection generates synthetic OOD data and trains MLPs to distinguish ID and OOD targets, with new evaluation protocols for realistic scenarios (Kösel et al., 2024).
  • Autonomous Driving. Autonomous driving is a critical application of OOD detection. Recent work uses the CARLA simulator (Dosovitskiy et al., 2017) to evaluate OOD detection performance in driving scenarios (Mao et al., 2024).  Mao et al. (2024) propose a language-augmented latent representation method, leveraging the image-text cosine similarity from the CLIP model to improve transparency and controllability in detection. Another recent work focus on real-time OOD perception in trajectory prediction, further enhancing vehicle safety (Banerjee et al., 2024). In real-world applications, Tesla’s “Shadow Mode” (tes, ) identifies cases where AI predictions differ from human drivers in real traffic—these OOD samples are collected and used to retrain and improve autonomous driving models.
  • Medical Image Analysis. OOD detection is vital in medical imaging, using datasets like CIFAR-10 and Kvasir-Capsul (Smedsrud et al., 2021) depending on the image category (Tan et al., 2024). Recent studies focus on reliable OOD detection methods for digital pathology (Oh et al., 2024) and test-time augmentation for improving OOD detection in gastrointestinal vision (Pokhrel et al., 2024). In practice, IDx-DR (Dis, ), the first FDA-approved autonomous AI diagnostic system, implements OOD detection by rejecting images that fail to meet quality standards or contain anomalies, deferring these cases to human experts for assessment.
  • Human Action Recognition. OOD detection plays a crucial role in ensuring the robustness of Human Action Recognition (HAR) models, particularly when faced with previously unseen actions. By leveraging advanced techniques such as attention-based debiasing (Sim et al., 2023), generative feature synthesis (Mandal et al., 2019) and energy-based skeleton modeling (Xu et al., 2024b), recent research has significantly improved OOD detection performance. These methods are evaluated on benchmark datasets like UCF-101 (Soomro et al., 2012), HMDB-51 (Kuehne et al., 2011) and Kinetics-400 (Kay et al., 2017).
  • Solar Image Analysis. In space weather forecasting, OOD detection serves as a crucial tool for identifying solar anomalies. Recent unsupervised  (Giger and Csillaghy, 2024; Das et al., 2024)approaches leverage Solar Dynamics Observatory data to enhance detection accuracy.
  • Facial Expression Recognition. In facial expression analysis, the challenge of distinguishing unknown expressions from known categories has led to advanced OOD detection methods. Recent approach (Zhang et al., 2024b) focuses on leveraging the unique characteristics of facial expression features, incorporating attention map consistency and cycle training mechanisms to effectively identify OOD samples.
  • Facial Authentication. In facial authentication systems, OOD detection is increasingly important for ensuring robust and secure identity verification. The FOOD framework (Kahya et al., 2024) exemplifies this trend by employing 60 GHz FMCW radar and a convolutional encoder-decoder architecture to achieve both accurate classification and reliable anomaly detection.
  • Earth Observation Imagery. For earth observation imagery, OOD detection enhances the identification of rare or unexpected events, such as natural disasters. Diffusion models like ODEED (Le Bellier and Audebert, 2024) have demonstrated strong performance in detecting OOD samples, with evaluations conducted on benchmark datasets such as SpaceNet 8 to validate their effectiveness.